OpenStudy (anonymous):

You have 15 pennies in your pocket. Of those pennies, 3 are Canadian. Suppose you pick a penny out of your pocket at random. Find probability (not canadian). A. 5 B. 1/5 C. 6/5 D. 4/5

OpenStudy (bahrom7893):

not canadian, therefore 15-3 = 12 non-canadian pennies thus: 12/15 = 4/5

OpenStudy (anonymous):

P ( canadian)= 3/15 = 1/5 -> P (not canadian) = 1 - 1/5 = 4/5